# On-Demand JSON: A Better Way to Parse Documents?

### v1

@@ Abstract
JSON is a popular standard for data interchange on the Internet. Ingesting JSON documents can be a performance bottleneck. A popular parsing strategy consists in converting the input text into a tree-based data structure—sometimes called a Document Object Model or DOM. We designed and implemented a novel JSON parsing interface—called On-Demand—that appears to the programmer like a conventional DOM-based approach. However, the underlying implementation is a pointer iterating through the content, only materializing the results (objects, arrays, strings, numbers) lazily. On recent commodity processors, an implementation of our approach provides superior performance in multiple benchmarks. To ensure reproducibility, our work is freely available as open source software. Several systems use On Demand: e.g., Apache Doris, the Node.js JavaScript runtime, Milvus, and Velox.
@@
"Width": 800,
"Height": 600,
"Title": "View from my room",
"Url": "http://ex.com/img.png",
"Private": false,
"Thumbnail": {
"Url": "http://ex.com/th.png",
"Height": 125,
"Width": 100
"array": [
116,
943,
234
"Owner": null
@@ 3 On-Demand Front-End
Conceptually, parsing with On Demand begins with the creation of a parser instance (ondemand::parser). The parser is responsible for memory allocation and it can be reused from document to document. Given a parser instance, the user calls the iterate method on a string input which returns a document instance (ondemand::document). The method calls the simdjson indexing routine, and writes an index in the parser instance. Afterward, the front-end consists essentially in a pointer over the document which we call internally a json_iterator: the iterator resides within the document instance.
@@
From the point of view of the programmer, we arx create arrays and objects (ondemand::array and ondemand::object). Yet these instances are mere thin wrappers and the creation of the instance translates into a check for the presence of the right structural characters (‘[’, ‘{’, ‘]’, ‘}’). They record their starting position and their document depth. They provide access to corresponding iterators over the elements of the arrays or the fields of the object.
@@
-array | ], "Owner": null } | 2 |
@@
Fig. 4 illustrates how the On Demand front-end might work with a non-trivial example. Given a JSON document made of an array of objects, we initiate a loop, accessing each element of the array as an object (car). Given such an object, we may query some fields (e.g., model) but not others (e.g., make): unqueried fields are skipped as much as possible. We may cast a value to a string, to an integer or to a floating-point value. In this particular example, if the value does not match the desired type, then a C++ exception would be thrown. We also support arbitrary schema through the type() method: it indicates whether the current node is potentially a number, a string, a Boolean, an array, an object, etc.
The simdjson library offers both a conventional DOM implementation and an On Demand approach. The On Demand approach requires less memory in general. While both strategies materialize the index, which uses one word (four bytes) per pseudo-structural character, the On Demand approach may not require any other intermediate memory.
In the DOM implementation of simdjson, a tree-like data structure is written during parsing, which is later accessed by the programmer. This data structure simply does not exist in the On Demand approach.
@@
The On Demand design has some known limitations. On Demand operates with a single pointer within the document. This means that if a user needs to process multiple components of the documents simultaneously, they must either materialize the data into their own data structure or rewind the iterator to earlier sections of the document. Rewinding the iterator may lead to reparsing the same JSON data multiple times, resulting in a performance penalty.
To illustrate this effect, consider a user who acquires an object instance and repeatedly queries a key value (e.g., object["mykey"]). Doing so might require scanning the JSON data of the object multiple times for the same key. In this scenario, the programmer should attempt to materialize the value once because On Demand does not currently support caching.
An anti-pattern for On Demand would be this code sequence:
@@
In general, On Demand provide an interface that appears like a DOM, but it has different performance characteristics and it requires some care to provide optimal performance. We therefore encourage the programmers to run their performance-sensitive code with logging enabled, so that the programmer can see the various tasks that the front-end has to carry.
The On-Demand provides complete error handling, and it is even possible to know where (in the JSON document) the error occurred, but it can be surprising for some users that an error is triggered after they have seemingly loaded the JSON document. Indeed, the On Demand front-end generally accepts invalid documents, and it only validates the components that the user parses, as they are being parsed. It requires slightly more care on the part of the programmer because the errors can happen at more locations in their code.
@@ 4 Experiments
Table 3 presents the processing speed for various tasks. We define the speed as the size of the input divided by the estimated elapsed time. Performance timings are heavily skewed toward the minimum; they do not follow a normal distribution [28]. Accordingly, we present the best processing speed (out of all runs) and the average processing speed. Except for the json2msgpack task, the On-Demand front-end is more than twice as fast as any competitor, except maybe for the DOM version of simdjson. Considering the geometric mean of the best results, we find that On Demand is 70% faster than the conventional (DOM-based) simdjson, it is over 2.5 times faster than yyjson, and nearly eight times faster than RapidJSON. On Demand is nearly 50 times faster than JSON for Modern C++.
task | simdjson (On Demand) | simdjson (DOM) | yyjson | RapidJSON | JSON for Modern C++ |
@@
Table 4 presents the number of CPU instructions retired by input byte. The number of instructions required is stable in our tests (within 1%). We find that the On-Demand front-end requires far fewer instructions than the competitors. It is indicative of a low number of mispredicted branches and of limited data dependencies. When considering the geometric mean, On Demand uses 60% of the instructions of the conventional (DOM-based) simdjson, half of yyjson’s instructions, and nearly eight times fewer instructions than RapidJSON. Compared to JSON for Modern C++, On Demand requires over 40 times fewer instructions. In general, On Demand is even faster than the numbers suggest, since it typically achieves a high number of instructions retired per unit of time. This is indicative of a low number of mispredicted branches and limited data dependencies.
Fig. 5 presents the speed and number of instructions in graphical form. JSON for Modern C++ is much slower than the alternatives and it is difficult to represent it graphically on the same scale: we omit its instruction counts in the plot, and its best speed is barely visible. We see graphically that the benefits of On Demand vary significantly depending on the task: they are modest for json2msgpack and much more significant for find tweet. We explain this difference by the fact that in json2msgpack, the entire document must be parsed into a data structure, and On Demand is thus unable to make gains due to its ability to skip some content. On the contrary, in the find tweet task, much of the JSON processing can be skipped without adverse consequence. On Demand still provides a benefit with the json2msgpack task because the data is written directly to the desired output without an intermediate storage.
task | simdjson (On Demand) | simdjson (DOM) | yyjson | RapidJSON | JSON for Modern C++ |
@@ 5 Conclusion
There are implementation design strategies that warrant further investigation. For example, our implementation relies on a precomputed index, giving us the location of structural characters: it makes our implementation convenient and the construction of the index is fast, but an On-Demand front-end may not require an index. Yet an even richer document index might prove more beneficial in some cases: instead of merely indexing the location of the objects, arrays and values, the index could be built with more information about the schema: e.g., the index could inform us that we have encountered an array made of integers. Future work should consider heterogeneous computing [29] with On Demand: the indexing could be computed by a GPU [15] or on an FPGA [12, 13, 14] and passed on to the On Demand front-end executed on general-purpose processor.

### v3

@@ Abstract
JSON is a popular standard for data interchange on the Internet. Ingesting JSON documents can be a performance bottleneck. A popular parsing strategy consists in converting the input text into a tree-based data structure—sometimes called a Document Object Model or DOM. We designed and implemented a novel JSON parsing interface—called On-Demand—that appears to the programmer like a conventional DOM-based approach. However, the underlying implementation is a pointer iterating through the content, only materializing the results (objects, arrays, strings, numbers) lazily. On recent commodity processors, an implementation of our approach provides superior performance in multiple benchmarks. To ensure reproducibility, our work is freely available as open source software. Several systems use On-Demand: e.g., Apache Doris, the Node.js JavaScript runtime, Milvus, and Velox.
@@
| ⬇ { "Width": 800, "Height": 600, "Title": "View from my room", "Url": "http://ex.com/img.png", "Private": false, "Thumbnail": { "Url": "http://ex.com/th.png", "Height": 125, "Width": 100 }, "array": [ 116, 943, 234 ], "Owner": null } |
@@ 3 On-Demand Front-End
Conceptually, parsing with On-Demand begins with the creation of a parser instance (ondemand::parser). The parser is responsible for memory allocation and it can be reused from document to document. Given a parser instance, the user calls the iterate method on a string input which returns a document instance (ondemand::document). The method calls the simdjson indexing routine, and writes an index in the parser instance. Afterward, the front-end consists essentially in a pointer over the document which we call internally a json_iterator: the iterator resides within the document instance.
@@
From the point of view of the programmer, we create arrays and objects (ondemand::array and ondemand::object). Yet these instances are mere thin wrappers and the creation of the instance translates into a check for the presence of the right structural characters (‘[’, ‘{’, ‘]’, ‘}’). They record their starting position and their document depth. They provide access to corresponding iterators over the elements of the arrays or the fields of the object.
@@
| -array | ], "Owner": null } | 2 |
@@
Fig. 4 illustrates how the On-Demand front-end might work with a non-trivial example. Given a JSON document made of an array of objects, we initiate a loop, accessing each element of the array as an object (car). Given such an object, we may query some fields (e.g., model) but not others (e.g., make): unqueried fields are skipped as much as possible. We may cast a value to a string, to an integer or to a floating-point value. In this particular example, if the value does not match the desired type, then a C++ exception would be thrown. We also support arbitrary schema through the type() method: it indicates whether the current node is potentially a number, a string, a Boolean, an array, an object, etc.
The simdjson library offers both a conventional DOM implementation and an On-Demand approach. The On-Demand approach requires less memory in general. While both strategies materialize the index, which uses one word (four bytes) per pseudo-structural character, the On-Demand approach may not require any other intermediate memory.
In the DOM implementation of simdjson, a tree-like data structure is written during parsing, which is later accessed by the programmer. This data structure simply does not exist in the On-Demand approach.
@@
The On-Demand design has some known limitations. On-Demand operates with a single pointer within the document. This means that if a user needs to process multiple components of the documents simultaneously, they must either materialize the data into their own data structure or rewind the iterator to earlier sections of the document. Rewinding the iterator may lead to reparsing the same JSON data multiple times, resulting in a performance penalty.
To illustrate this effect, consider a user who acquires an object instance and repeatedly queries a key value (e.g., object["mykey"]). Doing so might require scanning the JSON data of the object multiple times for the same key. In this scenario, the programmer should attempt to materialize the value once because On-Demand does not currently support caching.
An anti-pattern for On-Demand would be this code sequence:
@@
In general, On-Demand provide an interface that appears like a DOM, but it has different performance characteristics and it requires some care to provide optimal performance. We therefore encourage the programmers to run their performance-sensitive code with logging enabled, so that the programmer can see the various tasks that the front-end has to carry.
The On-Demand provides complete error handling, and it is even possible to know where (in the JSON document) the error occurred, but it can be surprising for some users that an error is triggered after they have seemingly loaded the JSON document. Indeed, the On-Demand front-end generally accepts invalid documents, and it only validates the components that the user parses, as they are being parsed. It requires slightly more care on the part of the programmer because the errors can happen at more locations in their code.
@@ 4 Experiments
Table 3 presents the processing speed for various tasks. We define the speed as the size of the input divided by the estimated elapsed time. Performance timings are heavily skewed toward the minimum; they do not follow a normal distribution [28]. Accordingly, we present the best processing speed (out of all runs) and the average processing speed. Except for the json2msgpack task, the On-Demand front-end is more than twice as fast as any competitor, except maybe for the DOM version of simdjson. Considering the geometric mean of the best results, we find that On-Demand is 70% faster than the conventional (DOM-based) simdjson, it is over 2.5 times faster than yyjson, and nearly eight times faster than RapidJSON. On-Demand is nearly 50 times faster than JSON for Modern C++.
| task | simdjson (On-Demand) | simdjson (DOM) | yyjson | RapidJSON | JSON for Modern C++ |
@@
Table 4 presents the number of CPU instructions retired by input byte. The number of instructions required is stable in our tests (within 1%). We find that the On-Demand front-end requires far fewer instructions than the competitors. It is indicative of a low number of mispredicted branches and of limited data dependencies. When considering the geometric mean, On-Demand uses 60% of the instructions of the conventional (DOM-based) simdjson, half of yyjson’s instructions, and nearly eight times fewer instructions than RapidJSON. Compared to JSON for Modern C++, On-Demand requires over 40 times fewer instructions. In general, On-Demand is even faster than the numbers suggest, since it typically achieves a high number of instructions retired per unit of time. This is indicative of a low number of mispredicted branches and limited data dependencies.
Fig. 5 presents the speed and number of instructions in graphical form. JSON for Modern C++ is much slower than the alternatives and it is difficult to represent it graphically on the same scale: we omit its instruction counts in the plot, and its best speed is barely visible. We see graphically that the benefits of On-Demand vary significantly depending on the task: they are modest for json2msgpack and much more significant for find tweet. We explain this difference by the fact that in json2msgpack, the entire document must be parsed into a data structure, and On-Demand is thus unable to make gains due to its ability to skip some content. On the contrary, in the find tweet task, much of the JSON processing can be skipped without adverse consequence. On-Demand still provides a benefit with the json2msgpack task because the data is written directly to the desired output without an intermediate storage.
| task | simdjson (On-Demand) | simdjson (DOM) | yyjson | RapidJSON | JSON for Modern C++ |
@@ 5 Conclusion
There are implementation design strategies that warrant further investigation. For example, our implementation relies on a precomputed index, giving us the location of structural characters: it makes our implementation convenient and the construction of the index is fast, but an On-Demand front-end may not require an index. Yet an even richer document index might prove more beneficial in some cases: instead of merely indexing the location of the objects, arrays and values, the index could be built with more information about the schema: e.g., the index could inform us that we have encountered an array made of integers. Future work should consider heterogeneous computing [29] with On-Demand: the indexing could be computed by a GPU [15] or on an FPGA [12, 13, 14] and passed on to the On-Demand front-end executed on general-purpose processor.